PREX Meeting Notes, Mar 4, 2010

Present: Bob Michaels, Ole Hansen, Mindy Jen, Luis Mercado,
         Paul Souder, Krishna Kumar, Megan Friend, Gregg Franklin,
         Brian Quinn, Richard Holmes, Sasha Glamazdin, Paul King,
         Alexandre Camsonne, Zafar Ahmed, Diancheng Wang,
         Xiaoyan Deng, Xiaochao Zheng, Jon Wexler

GEMs -- Ole, KK, Alex

  Krishna: mechanical parts in shop, assembly next Wed, Thurs.
  Ole:  Only 6 chambers available (would need 8).  Enough 
             for 1 arm for sure, can make do.
        Two Italian GEMs in EEL.
        Italian collaborators arriving today.
        Guido, 1 grad student, 1 tech (Bob note: sorry I didn't
                                        catch their names)
        Nilanga -- bringing 2 more GEMs Monday, +2 by end
             of next week.
        Have all cables, low-V supplies.  Some question
             if we have HV supplies.
   Alex: Setting up Gassiplex DAQ readout.
         Timing looks good.

Moller -- Sasha

   Major problem: vacuum leaks !
   No time estimate yet for when the leaks will be fixed.
   Fixed MPS signal shape.
   Doing cosmics run checkout.
   Working with MCC: there is (1) new Moller FSD and 
       (2) MCC will test the "user" mode [user mode is a 
       pulsed beam to hall A while CW to hall B -- the idea 
       is to reduce target heating while being non-invasive 
       to hall B.]
   Sasha will present new Moller procedure to MCC ops.

Injector Beam Studies --

   Kent, Rupesh, Mark, and Gordon didn't make it to the meeting, 
   but Bob can report that they, together with Riad, have been
   extremely busy and productive this past ~1.5 weeks with 
   studies of the polarized source setup and beam at the injector.

DAQ -- Luis

   Plan is to finish the DAQ during Fri - Sun (Mar 5 - 7).
   Install full complement of ADCs.
   Swap out one bad ADC.
   Continue PMT linearity tests.
   Move DAQ to adaql5 to solve the 240 Hz / ET problem
                            (discussed in HALOG last few days)

Mindy -- Compton

   Finished pulser tests of CompSimple (see Compton Elog).
   Modified CMU code to incorporate Electron detector
     and to find coincidences for CompSimple.
     -- checking HAPPEX-III data to verify code still works
     -- checking standalone ETROC data
     -- looks good so far
Zafar -- Cavities

   Reproduced the plots from previous low-current running.
   Analysis code is ready.
   Prepared the amplifiers for low-current running.
   Checking connections to DAQ, running DAQ to verify 
       the signals work.
   Working with John Musson to learn all procedures.

Megan -- Compton

   LED pulser fixed.
   Checked Linearity of PMTs for PREX energy range.
   Looks good, has 2 PMTs ready.
   See HAPLOG for details.

Richard Holmes -- Compton

   See his recent HAPLOG entries.
   Monte Carlo vs data for response function during HAPPEX-3.
   Polarization depends on cuts -- anomolous, not understood.
   Possible offset in energy scale, or other problem
   Gregg: speculates shift may be caused by pileup.
         Discussion about that ensues.
         (Bob: check rate dependence ?)
   Issues not yet resolved.

 Next meeting:   Thurs Mar 11 at 3 pm

Index of Meetings